A tobacco company advertises that the average nicotine content of its cigarettes is at

most 14 milligrams. A consumer protection agency wants to determine whether the

average nicotine content is in fact greater than 14. A random sample of 300 cigarettes

of the company’s brand yield an average nicotine content of 14.6 and a standard

deviation of 3.8 milligrams. Determine the level of significance of the statistical test of

the agency’s claim that 𝜇 is greater than 14. If 𝛼 = 0.01, is there significant evidence

that the agency’s claim has been supported by the data?


1
Expert's answer
2021-12-01T13:32:52-0500

The following null and alternative hypotheses need to be tested:

"H_0:\mu\leq 14"

"H_1:\mu>14"

This corresponds to a right-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.01," "df=n-1=299" degrees of freedom, and the critical value for a right-tailed test is "t_c =2.338884."

The rejection region for this right-tailed test is "R = \{t: t > 2.338884.\}"

The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{14.6-14}{3.8/\sqrt{300}}"

"\approx 2.734817"

Since it is observed that "t = 2.734817>2.338884 = t_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach: The p-value for right-tailed, "\alpha=0.01, df=299" degrees of freedom, "t=2.734817" is"p=0.003307," and since "p=0.003307<0.01=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu"

is greater than 14, at the "\alpha = 0.01" significance level.